# From Trace to Line: LLM Agent for Real-World OSS Vulnerability Localization

### v1

@@ Abstract
Large language models show promise for vulnerability discovery, yet prevailing methods inspect code in isolation, struggle with long contexts, and focus on coarse function or file level detections—offering limited actionable guidance to engineers who need precise line-level localization and targeted patches in real-world software development. We present T2L-Agent (Trace-to-Line Agent), a project-level, end-to-end framework that plans its own analysis and progressively narrows scope from modules to exact vulnerable lines. T2L-Agent couples multi-round feedback with an Agentic Trace Analyzer (ATA) that fuses runtime evidence—crash points, stack traces, and coverage deltas with AST-based code chunking, enabling iterative refinement beyond single pass predictions and translating symptoms into actionable, line-level diagnoses. To benchmark line-level vulnerability discovery, we introduce T2L-ARVO, a diverse, expert-verified 50-case benchmark spanning five crash families and real-world projects. T2L-ARVO is specifically designed to support both coarse-grained detection and fine-grained localization, enabling rigorous evaluation of systems that aim to move beyond file-level predictions. On T2L-ARVO, T2L-Agent achieves up to 58.0% detection and 54.8% line-level localization, substantially outperforming baselines. Together, the framework and benchmark push LLM-based vulnerability detection from coarse identification toward deployable, robust, precision diagnostics that reduce noise and accelerate patching in open-source software workflows. Our framework and benchmark are publicly available as open source at https://github.com/haoranxi/T2LAgent.
@@ 1 Introduction
Software vulnerabilities now occur at unprecedented scale and cost. In 2023, more than 29,000 Common Vulnerabilities and Exposures (CVEs) were recorded CVE Details (2024). In the first half of 2025, 1,732 data breaches were reported—an 11% increase Indusface (2024). The economic toll is mounting: software supply-chain attacks are projected to cost the global economy 80.6 billion annually by 2026, up from 45.8 billion in 2023 Dark Reading (2024); some estimates put total damages at \$9.5 trillion in 2024 Liu et al. (2024). Critically, 14% of breaches in 2024 began with vulnerability exploitation—nearly triple the prior year. Despite advances in automated vulnerability detection, effectiveness in localization remains underexplored Zhang et al. (2024b), leaving developers with coarse, file-level predictions while 32% of critical vulnerabilities remain unpatched for over 180 days. These vulnerabilities are exploding in volume and growing ever more severe, driving escalating security and economic risks worldwide.
Motivation. The rise of Large Language Models (LLMs) has accelerated AI-driven automation across software development and cybersecurity Zhang et al. (2025). From code assistants like GitHub, Inc. (2021) to automated security analysis and automation Nunez et al. (2024). LLMs show strong aptitude for understanding complex codebases and flagging potential issues Divakaran & Peddinti (2024). This momentum naturally extends to vulnerability detection, spanning model-level approaches such as fine-tuning LLMs for classification—and agentic frameworks that harness LLM reasoning for automated analysis tmylla and contributors (2024).
Yet today’s LLM-based detection methods face practical barriers. Most operate at the function level, asking whether a fragment is vulnerable rather than pinpointing where the flaw lies Zhang et al. (2024b), Sovrano et al. (2025), yielding guidance too coarse for real remediation. Evaluations also rely heavily on lightweight, synthetic datasets that miss the complexity of production systems Guo et al. (2024).
@@
Contribution. This work makes three contributions: (1) Task formulation. We frame project-level vulnerability discovery as two structured tasks: chunk-level detection and line-level localization. AST-based chunking adapts large codebases to LLM context while preserving semantics, enabling refinement from coarse predictions to exact lines,bridging research setups and real-world needs for LLMs. (2) T2L-ARVO benchmark. We present the first benchmark for agentic fine-grained localization, featuring 50 expert verified cases across five vulnerability types with balanced category distribution. T2L-ARVO enables realistic, project-scale evaluation of LLM-based systems. (3) T2L-Agent framework. We propose a multi-agent system with the following innovation: (i) a Trace Analyzer (Agentic Trace Analyzer) integrating the ensemble of tools such as static analysis, sanitizers, and runtime monitoring for observability as human debuggers; (ii) a Proposal module (Divergence Tracing) that iteratively forms and tests vulnerability hypotheses; and (iii) a two-stage pipeline (Detection Refinement) that first detects coarse chunks and then verifies exact lines with runtime feedback.
T2L-Agent achieves up to 58% chunk-level detection and 54.8% exact line localization accuracy on T2L-ARVO, significantly reducing developer effort. Its iterative, feedback-driven workflow mirrors how engineers debug—combining reasoning with runtime evidence—while scaling effectively to large, real-world codebases through a lightweight agent architecture.
@@ 2 Background
Vulnerability Localization. Classical localization combines static and dynamic analyses with retrieval and graph methods. Static slicing selects statements that may affect a slicing criterion to form behavior-preserving slices for debugging and comprehension Weiser (1981). Dynamic slicing refines this by deriving input-specific slices from execution-time dependence graphs, yielding higher precision near crashes or variables Agrawal & Horgan (1990). Information-retrieval approaches treat bug reports as queries, ranking files/methods by textual and historical signals to surface likely locations Zhou et al. (2012); Saha et al. (2013); Wang & Lo (2014). Code Property Graphs unify classic program-analysis concepts in a joint representation, enabling scalable pattern searches that uncovered 18 previously unknown vulnerabilities Yamaguchi et al. (2014).
AI for Cybersecurity. Recent agentic systems couple LLM planning with tool execution and feedback in runnable environments. Planner–Executor frameworks (e.g., D-CIPHER) generate detailed plans and dispatch executors, closing the loop with dynamic feedback and reporting state-of-the-art CTF results Udeshi et al. (2025); Shao et al. (2024); Yang et al. (2023); Zhang et al. (2024a). CRAKEN adds knowledge-based execution via a security knowledge base and retrieval pipeline, improving performance on NYU CTF Bench over prior baselines Shao et al. (2025). EnIGMA shows that interactive tool use helps discover and exploit vulnerabilities, outperforming purely text-only setups across multiple CTFs Abramovich et al. (2025). Pen-testing systems (PentestGPT, PentestAgent) coordinate specialized modules or multi-agent collaboration to automate vulnerability analysis and exploitation in realistic scenarios Deng et al. (2024); Shen et al. (2025).
LLM Agentic Systems. LLMs increasingly automate scientific workflows across chemistry, quantum, and biology Bran et al. (2023); Basit et al. (2025); Jin et al. (2024). ChemCrow augments LLMs with 18 expert-designed tools to execute multi-step syntheses and discovery tasks end-to-end Bran et al. (2023); ChemGraph integrates graph neural networks and simulators to drive computational chemistry pipelines Pham et al. (2025); MDCrow orchestrates chain-of-thought with >40>40 tools for file handling, simulation setup, and analysis Campbell et al. (2025). In life sciences, ProtAgents uses multi-agent collaboration for protein discovery Ghafarollahi & Buehler (2024), while VibeGen introduces a dual-model design guided by normal-mode vibrations and assessed with a protein predictor Ni & Buehler (2025). Collectively, these systems reduce manual configuration and broaden access via natural-language interfaces.
@@ 3 Related Work
For vulnerability localization task, LLM also shows great potential. Early learning based vulnerability detection papers improve localization accuracy by training models to classify whether a unit is vulnerable. LLMAOYang et al. (2024a) fine-tunes LLMs on small, manually curated buggy programs, while BAPStein et al. (2025) learns state-of-the-art vulnerability localization without any direct localization labels, outperforming traditional baseline over eight benchmarks. However, following analysis also found data issues in widely used datasets such as Big-Vul and DevignZhou et al. (2019), making people doubt about performance numbers and highlighting the needs for a realistic evaluation settingsCroft et al. (2023).
@@
LLMAO 2024a
BAP 2025
GenLoc 2025
AgentFL 2024
CoSIL 2025
AutoFL 2024
LineVul 2022
LOVA 2024
MatsVD 2024
xLoc 2024b
LLM4FL 2024
MemFL 2025
@@
Many prior works frame vulnerability localization at the file or function level, but this coarse granularity often fails to provide actionable guidance for developers. GenLoc Asad et al. (2025) identifies potentially vulnerable files from bug reports and iteratively analyzes them using code exploration tools. AgentFL Qin et al. (2024) applies a multi-agent framework for function-level localization, modeling the task as a three-step pipeline with specialized agents and tools. CoSIL Jiang et al. (2025) narrows the function-level search space using module call graphs and iteratively traverses them for relevant context. Similarly, AutoFL Kang et al. (2024) prompts LLMs to localize method-level vulnerabilities via function-call navigation, showing that multi-step reasoning helps overcome context length limits.
To offer more precise guidance, recent studies have shifted toward line- or statement-level localization. LineVul Fu & Tantithamthavorn (2022) uses a Transformer-based classifier for line-level prediction, while LOVA Li et al. (2024) introduces a self-attention framework to score and highlight vulnerable lines. MatsVD Weng et al. (2024) enhances statement-level localization using dependency-aware attention, and xLoc Yang et al. (2024b) learns multilingual, task-specific knowledge for bug detection and localization.
Building on these efforts, LLM4FL Rafi et al. (2024) proposes a multi-agent framework leveraging graph-based retrieval and navigation to reason about failure causes. MemFL Yeo et al. (2025) introduces external memory to incorporate project-specific knowledge, improving localization in complex, repository-scale systems.
Collectively, these works push localization from file to line level and increasingly adopt multi-agent strategies for subtask coordination. However, most still rely on limited runtime evidence, single-pass predictions, or benchmarks that lack realistic project settings. Our T2L-Agent addresses these limitations with a planner-executor framework that incorporates runtime signals, enables dynamic feedback loops, and achieves line-level localization in real-world repository environments.
@@ 4 Methods
4.1 T2L Framework
T2L-Agent uses a hierarchical planner–executor architecture Udeshi et al. (2025) that breaks vulnerability localization into evidence collection, hypothesis generation, and iterative refinement. Unlike single-pass static analyzers, it follows a human-like workflow: gather runtime signals, correlate them with code structure, and progressively shrink the search space.
@@ 4.1 T2L Framework
Evidence Tracing T2L Planner coordinates repository analysis and runtime evidence capture in a single, structured pipeline. First, code-structure analysis partitions the codebase into function-aligned, semantically coherent units that preserve syntactic relationships while fitting LLM context windows; known patch locations are also indexed to establish evaluation baselines. Next, runtime evidence becomes the cornerstone: Sanitizer records memory-violation patterns, allocation traces, and stack frames, while interactive debugging (GDB/LLDB) provides symbolic context via backtraces and variable-state snapshots at crash points. This dual-layer design yields comprehensive, observable crash logs rather than speculative static signals. The integrated workflow is a key T2L innovation; we detail it in Sec. 4.2.
Two-stage Refinement T2L Executor Agent refines candidates iteratively. It bootstraps broad candidate regions from crash logs, using LLM code comprehension to connect symptoms to likely causes, and emits ranked candidate lists with confidence level and rationale. Subsequent passes run targeted source inspections, checking for patterns such as missing bounds checks, uninitialized variables, and improper memory management, then produce new or improved candidates from the inspected code and validate against ground-truth patches.
@@
4.2 Agentic Trace Analyzer (ATA)
The Agentic Trace Analyzer bridges static code analysis and runtime behavior—long a blind spot in vulnerability-detection pipelines. Most systems analyze code in isolation and miss the symptom-to-cause chain. We close this gap with an end-to-end, multi-source evidence pipeline: targets run in Docker for reproducible, consistent environments; Tree-sitter partitions repositories into semantically meaningful chunks that preserve structure for precise slice extraction; executions are instrumented with analysis toolkit such as Sanitizers, Debuggers and Static Analyzers to capture stack traces, memory-violation reports, register states, and control-flow cues. Details of the ATA tool list that T2L originally supports are provided in  A.5.
We apply a hierarchical refinement: start from coarse crash signatures and narrow to concrete code locations. A dynamic evidence graph correlates runtime observations with static features for cross-validation, while we score and sort candidates by how well they match across multiple signals—syntax, semantics, and execution traces (syntactic patterns, semantic embeddings, execution-trace alignment). Mirroring how engineers debug—alternating reading with running—the analyzer in T2L-Agent seeds initial candidates through static–dynamic correlation, then iteratively refines them against real source slices in feedback loops. The introduction of ATA brings fewer single-shot failures, improved compute efficiency, and behavior-anchored decisions that enable precise, line-level localization even in large, tightly coupled codebases. We disable the ATA and allow LLM to localize vulnerabilities, an example is shown in Figure 4 (c). Without ATA, LLM could not successfully localize vulnerabilities compared to (a) and (b) with ATA enabled.
4.3 Fine-grained detection
Divergence Tracing. Recognizing that complex vulnerabilities often involves multiple files and functions, T2L-Agent also uses divergence tracing to explore multiple hypotheses in parallel from the same crash signature. This feature is inspired by how modern LLM interfaces offer multiple response variations for users, using the LLM’s variability to ensure comprehensive coverage of potential vulnerability locations. Rather than committing to a single chain of thought, it expands several in parallel and returns a ranked list of candidate sites across the search space. This surfaces correct localizations that were not top ranked initially and is especially helpful for bugs that span multiple modules. From Figure 4 (b), we can observe that the divergence tracing generates more localization candidates and matched more vulnerable lines in this round.
Detection Refinement. The detection refinement process begins based on crash logs and initial localization candidate regions at the first step. Rather than exhaustively examining all potentially relevant code, the agent selects source code slices based on crash signatures, stack trace information, and vulnerability patterns. On a second pass, it rereads those slices to find missed patterns by checking syntactic cues, semantic links such as data and control flow, and alignment with runtime evidence. The refinement process operates iteratively to help the agent correct early mistakes and discover vulnerabilities that are not immediately obvious from crash logs alone, particularly for complex vulnerabilities involving memory corruption where the crash point may be far away from the actual vulnerability. As Figure 4 (a) shows, based on the source code LLM interested, the refinement process successfully locates new lines that were not found during first step with only the runtime evidence.

### v2

@@ Abstract
Large language models show promise for vulnerability discovery, yet prevailing methods inspect code in isolation, struggle with long contexts, and focus on coarse function or file level detections which offers limited actionable guidance to engineers who need precise line-level localization and targeted patches in real-world software development. We present T2L-Agent (Trace-to-Line Agent), a project-level, end-to-end framework that plans its own analysis and progressively narrows scope from modules to exact vulnerable lines. T2L-Agent couples multi-round feedback with an Agentic Trace Analyzer (ATA) that fuses runtime evidence such as crash points, stack traces, and coverage deltas with AST-based code chunking, enabling iterative refinement beyond single pass predictions and translating symptoms into actionable, line-level diagnoses. To benchmark line-level vulnerability discovery, we introduce T2L-ARVO, a diverse, expert-verified 50-case benchmark spanning five crash families and real-world projects. T2L-ARVO is specifically designed to support both coarse-grained detection and fine-grained localization, enabling rigorous evaluation of systems that aim to move beyond file-level predictions. On T2L-ARVO, T2L-Agent achieves up to 58.0% detection and 54.8% line-level localization, substantially outperforming baselines. Together, the framework and benchmark push LLM-based vulnerability detection from coarse identification toward deployable, robust, precision diagnostics that reduce noise and accelerate patching in open-source software workflows. Our framework and benchmark are publicly available as open source at https://github.com/haoranxi/T2LAgent.
@@ 1 Introduction
Software vulnerabilities now occur at unprecedented scale and cost. In 2023, more than 29,000 Common Vulnerabilities and Exposures (CVEs) were recorded cve_details_2024. In the first half of 2025, 1,732 data breaches were reported–an 11% increase indusface_2024. The economic toll is mounting: software supply-chain attacks are projected to cost the global economy 80.6 billion annually by 2026, up from 45.8 billion in 2023 darkreading_supply_chain; some estimates put total damages at \$9.5 trillion in 2024 liu_makridis_galinkin_2024. Critically, 14% of breaches in 2024 began with vulnerability exploitation—nearly triple the prior year. Despite advances in automated vulnerability detection, effectiveness in localization remains underexplored zhang2024empirical, leaving developers with coarse, file-level predictions while 32% of critical vulnerabilities remain unpatched for over 180 days. These vulnerabilities are exploding in volume and growing ever more severe, driving escalating security and economic risks worldwide.
Motivation. The rise of Large Language Models (LLMs) has accelerated AI-driven automation across software development and cybersecurity zhang_et_al_2025. From code assistants like github_copilot_2021 to automated security analysis and automation nunez2024autosafecoder. LLMs show strong aptitude for understanding complex codebases and flagging potential issues divakaran2024llms. This momentum naturally extends to vulnerability detection, spanning model-level approaches such as fine-tuning LLMs for classification—and agentic frameworks that harness LLM reasoning for automated analysis awesome_llm4cybersecurity_2024.
Yet today’s LLM-based detection methods face practical barriers. Most operate at the function level, asking whether a fragment is vulnerable rather than pinpointing where the flaw lies zhang2024empirical, 10.1145/3715758. These tasks are also defined purely on static code, ignoring binaries, runtime behavior, and project-level diffs, further limiting their usefulness for real remediation. Evaluations also rely heavily on lightweight, synthetic datasets that miss the complexity of production systems springer_vulnerability_analysis.
@@
Contribution. This work makes three contributions: (1) New Task formulation. We define a new task for agentic AI, a runtime trace guided, project-level vulnerability detection, and formulate it as two structured subtasks: chunk-level detection and line-level localization. AST-based chunking adapts large codebases to LLM context while preserving semantics, enabling refinement from coarse predictions to exact lines, bridging research setups and real-world needs for LLMs. (2) T2L-ARVO benchmark. We present the first benchmark for agentic fine-grained localization, featuring 50 expert verified cases across five vulnerability types with balanced category distribution. T2L-ARVO enables realistic, project-scale evaluation of LLM-based systems. (3) T2L-Agent framework. We propose a multi-agent system with the following innovation: (i) a Trace Analyzer (Agentic Trace Analyzer) integrating the ensemble of tools such as static analysis, sanitizers, and runtime monitoring for observability as human debuggers; (ii) a Proposal module (Divergence Tracing) that iteratively forms and tests vulnerability hypotheses; and (iii) a two-stage pipeline (Detection Refinement) that first detects coarse chunks and then verifies exact lines with runtime feedback.
Overall, T2L-Agent provides a comprehensive baseline for this new vulnerability localization task. Unlike prior snippet-level detectors, it is designed to operate on fully reproducible real-world projects, using crashes, sanitizer reports, and project diffs as runtime evidence. On T2L-ARVO, it achieves up to 58% chunk-level detection and 54.8% exact line localization, showing that this runtime-evidence–guided setting is challenging yet tractable for LLM agents.
@@ 2 Background
Vulnerability Localization. Classical localization combines static and dynamic analyses with retrieval and graph methods. Static slicing selects statements that may affect a slicing criterion to form behavior-preserving slices for debugging and comprehension 10.5555/800078.802557. Dynamic slicing refines this by deriving input-specific slices from execution-time dependence graphs, yielding higher precision near crashes or variables agrawal1990dynamic. Information-retrieval approaches treat bug reports as queries, ranking files/methods by textual and historical signals to surface likely locations zhou2012should; 6693093; 10.1145/2597008.2597148. Code Property Graphs unify classic program-analysis concepts in a joint representation, enabling scalable pattern searches that uncovered 18 previously unknown vulnerabilities yamaguchi2014modeling.
AI for Cybersecurity. AI is widely applied to cybersecurity tasks with tool execution and feedback in runnable environments yang2023intercode. One prominent use is offensive security shao2024empirical, where agents reason about attack chains and interact with benchmark environments shao2024nyu; zhang2024cybench to simulate human CTF players udeshi2025d; shao2025craken; abramovichenigma. In parallel, pen-testing systems orchestrate modules or multi-agent collaboration to automate reconnaissance, vulnerability analysis, and exploitation against targets deng2024pentestgpt; shen2025pentestagent. Further work builds agent-based patching frameworks that localize vulnerabilities and synthesize fixes yu2025patchagent; xue2025pagent, and explores LLM-augmented red and blue teaming, where agents emulate adaptive attackers and assist defenders in threat hunting and incident investigation abuadbba2025promise; liu2025benchmarking.
LLM Agentic Systems. Beyond cybersecurity, large language models (LLMs) are increasingly applied to automate workflows across diverse domains shao2024survey, including AI for scientific research bran2023chemcrow; basit2025pennylang; jin2024prollm, software engineering yang2024swe, healthcare isern2016systematic, and hardware design wang2024llms. Many systems integrate long-term memory to retain intermediate results, environment states, and user preferences across sessions zhang2025survey. LLM as judge are often used to evaluate actions, select among tool outputs, and provide feedback for self-improvement  li2025generation. Some agents employ self-reflection and critique loops to iteratively refine their reasoning or code before execution renze2024self. Multi-agent and role-specialized architectures further enable collaboration among agents with distinct skills to solve complex, interleaved subtasks in parallel li2024survey.
@@ 3 Related Work
LLMs show strong potential on vulnerability localization tasks. Early learning-based detectors improve localization by training models to classify whether a unit is vulnerable. LLMAOyang2024large fine-tunes LLMs on small, manually curated buggy programs, while BAPstein2025s learns state-of-the-art vulnerability localization without any direct localization labels, outperforming traditional baseline over eight benchmarks. However, subsequent analyses revealed data issues in widely used datasets such as Big-Vul and Devignzhou2019devign, making people doubt about performance numbers and highlighting the need for realistic evaluation settings croft2023data.
@@
LLMAO yang2024large
BAP stein2025s
GenLoc asad2025leveraging
AgentFL qin2024agentfl
CoSIL jiang2025cosil
AutoFL kang2024quantitative
LineVul 9796256
LOVA li2024attention
MatsVD 10.1145/3671016.3674807
xLoc 10.1145/3660804
LLM4FL rafi2024multi
MemFL yeo2025improving
@@
Many prior works frame vulnerability localization at the file or function level, but this coarse granularity often fails to offers limited guidance for developers. GenLoc asad2025leveraging identifies potentially vulnerable files from bug reports and iteratively analyzes them using code exploration tools. AgentFL qin2024agentfl applies a multi-agent framework for function-level localization, modeling the task as a three-step pipeline with specialized agents and tools. CoSIL jiang2025cosil narrows the function-level search space using module call graphs and iteratively traverses them for relevant context. Similarly, AutoFL kang2024quantitative prompts LLMs to localize method-level vulnerabilities via function-call navigation, showing that multi-step reasoning mitigates context limits.
Recent studies have shifted toward line or statement level localization. LineVul 9796256 uses a Transformer-based classifier for line-level prediction, while LOVA li2024attention introduces a self-attention framework to score and highlight vulnerable lines. MatsVD 10.1145/3671016.3674807 enhances statement level localization with dependency-aware attention, and xLoc 10.1145/3660804 learns multilingual, task-specific knowledge for bug detection and localization. Building on these efforts, LLM4FL rafi2024multi proposes a multi-agent framework leveraging graph-based retrieval and navigation to reason about failure causes. MemFL yeo2025improving introduces external memory to incorporate project-specific knowledge, improving localization in complex, repository-scale systems. CThey push localization to line level and use more multi-agent coordination. However, most still rely on limited runtime evidence, single-pass predictions, or benchmarks that lack realistic project settings.
T2L formalizes a new task of runtime-trace-guided vulnerability localization at the project level. Prior learning-based approaches such as BAP, LLMAO, and LineVul rely on static code analysis, training models to classify whether a given code snippet is vulnerable based solely on syntactic and semantic patterns. These methods typically operate on isolated single files and produce coarse-grained predictions at the function or file level without leveraging dynamic execution information. In contrast, our task targets multi-file codebases and incorporates runtime evidence including reproducible crash logs, stack traces, and sanitizer reports to pinpoint exact vulnerable lines. This formulation more closely reflects real-world vulnerability triage workflows where developers rely on both static code inspection and dynamic program behavior to diagnose security flaws.
@@ 4 Methods
4.1 T2L Agent
T2L-Agent uses a hierarchical planner–executor design udeshi2025d that splits localization into evidence gathering, hypothesis generation, and iterative refinement.Unlike single-pass analyzers, it uses a human-like process: gather runtime signals, link them to code, and narrow the search space. Details of the tool list that T2L originally supports are provided in  A.5.
@@ 4.1 T2L Agent
Evidence Tracing T2L Planner coordinates repository analysis and runtime evidence capture in a single, structured pipeline. First, code-structure analysis partitions the codebase into function-aligned, semantically coherent units that preserve syntactic relationships while fitting LLM context windows; known patch locations are also indexed to establish evaluation baselines. Next, runtime evidence becomes the cornerstone: Sanitizer records memory-violation patterns, allocation traces, and stack frames, while interactive debugging (GDB/LLDB) provides symbolic context via backtraces and variable-state snapshots at crash points. This dual-layer design yields comprehensive, observable crash logs rather than speculative static signals. We collect multiple types of runtime and static evidence, including crash logs, sanitizer reports, stack traces, and static analyzer results, all evidence is merged into a single unified textual block and fed to the LLM at once. This design ensures that the LLM forms its reasoning based on the global evidence context, rather than aggregating per-evidence predictions to avoids issues of cross-signal inconsistency. The integrated workflow is a key T2L innovation; we detail it in Sec. 3.
Hypothesis Generation A “hypothesis” an LLM-generated candidate vulnerability produced from the unified evidence block of an existing vulnerable project in T2L system. The LLM proposes a list of file:line candidate pairs ranked by confidence. Each hypothesis is independently checked against the ground-truth patch diff, and candidates do not affect one another during evaluation.
Two-stage Refinement To ensure accurate detection, T2L uses a two-stage refinement process. It begins with coarse-grained filtering: the system derives broad candidate regions from crash logs, using LLM code comprehension to link symptoms to likely causes and produce a ranked list of candidates with confidence and rationale. It then identifies code fragments the LLM flagged as relevant in this first pass–such as missing bounds checks, uninitialized variables, or improper memory management. Next comes fine-grained refinement: these extracted fragments are added to the evidence for a second LLM pass, which updates the candidate set. Any newly generated locations are merged into the existing candidates. This refinement loop continues until the LLM produces no additional candidates or the cost budget is reached.
@@
Agentic Trace Analyzer (ATA) The Agentic Trace Analyzer (ATA) is the core module of T2L-Agent. Its purpose is to connect runtime behavior with relevant source code regions, addressing a gap in prior work that relies mainly on static signals. ATA constructs a unified “evidence block” by running targets in reproducible Docker, partitioning repositories into semantically meaningful chunks using tree-sitter, and instrumenting executions with analysis toolkit such as Sanitizers, Debuggers, and Static analyzers to collect stack traces, memory errors, and other runtime cues. The output is a unified “evidence block” which is then passed to T2L-Agent’s reasoning modules.
We use hierarchical refinement, starting from crash logs and narrowing to specific code. The agent cross-validates runtime signals with static features and ranks candidates by LLM confidence. Like human debugging, T2L-Agent seeds candidates via static–dynamic correlation and iteratively refines them with source–feedback loops. ATA brings fewer single-shot failures, improved compute efficiency, and behavior-anchored decisions that enable precise, line-level localization even in large, tightly coupled codebases. With ATA disabled, the LLM localizes on its own, as shown in Figure 4(c). Without ATA, the LLM fails to localize vulnerabilities compared with (a) and (b).
@@
4.2 T2L Features
Divergence Tracing. Recognizing that complex vulnerabilities often involves multiple files and functions, T2L-Agent also uses divergence tracing to explore multiple hypotheses in parallel from the same crash log. This feature is inspired by how modern LLM interfaces offer multiple response variations for users, using the LLM’s variability to ensure comprehensive coverage of potential vulnerability locations. Rather than committing to a single chain of thought, it expands several in parallel and returns a ranked list of candidate sites across the search space. This surfaces correct localizations that were not top ranked initially and is especially helpful for bugs that span multiple modules. From Figure 4 (b), we can observe that the divergence tracing generates more localization candidates and matched more vulnerable lines in this round.
Detection Refinement. The detection refinement begins based on crash logs and initial localization candidates. Each LLM-generated candidate gets a confidence label from very unlikely to very likely. We use these labels to sort and rank the candidates during the coarse-grained filtering. T2L agent selects code slices based on crash logs, stack trace information, and vulnerability patterns instead of exhaustively examining the full code snippet. In fine-grained refinement, we extract these slices and feed them back to the LLM along with the unified evidence, providing richer context surrounding the initially flagged locations.
The refinement process operates iteratively to help the agent correct early mistakes and discover vulnerabilities that are not immediately obvious from traces, particularly for complex vulnerabilities involving memory corruption where the crash point may be far away from the actual vulnerability. As Figure 4 (a) shows, based on the source code LLM interested, the refinement process successfully locates new lines that were not found during first step with only the runtime evidence.